# The Web unpacked: a quantitative analysis of global Web usage

### v1

@@ 1. Introduction
The importance of the Web, its lack of explicit management, and its vast potential have raised questions on what are its realized, actual characteristics in terms of main uses, applications, and content types, how visits are distributed among websites, who are the content publishers, how is the content typically produced, and much more. In the past, the Web has been thoroughly analyzed in terms of its content and graph properties (e.g., the number of webpages covering a given topic (Chakrabarti et al., 2002; Huizingh, 2000) and the number of hyperlinks connecting them (Adamic and Huberman, 2000b; Huberman, 2001)). However, previous accounts about the characteristics of the Web have primarily been based on metrics other than usage (e.g., link structure, number of websites, or companies’ market value) or were built on anecdotal accounts and non-quantitative data. Some exceptions are: (Adamic and Huberman, 2000a; Webster and Lin, 2002) that showed for a restricted subset of users and websites that visitors are distributed among websites in a power-law-like fashion; and (Agarwal and Sastry, 2022), that used Alexa’s ranking list to select the 100 most visited domains, Google trends to estimate their popularity and Fortiguard’s classification to verify the popularity of different topics, among other things.
@@ 2. Data Set
SimilarWeb, a company that combines all three methods, reportedly tracks a representative panel of users to ensure that all domains visited a minimum number of times are monitored (SimilarWeb, 2024b).
@@
The quality of SimilarWeb data has been evaluated in previous academic works and is considered good enough for relative ballpark measurements between domains. Jansen et al. used Google Analytics for 86 websites as a truth table and identified a systematic bias of 20% in SimilarWeb’s monthly visits estimates (Jansen et al., 2022). Prantl compared monthly visits from SimilarWeb and NetMonitor [^1] https://www.spir.cz/projekty/netmonitor/ for 485 Czech websites and found an average absolute difference of about 42% between these two sets of measurements (Prantl and Prantl, 2018). Enterprise reports have indicated similar scenarios (Hardwick, 2018; Diachuk et al., 2021). However, given that our data on monthly visits cover over seven orders of magnitude (from 5 thousand to 86 billion), the reported deviations do not significantly affect the overall distribution of visits among websites. The data still provides precise ranks, with a Pearson correlation of 95% between monthly visit estimates from SimilarWeb and Google Analytics (Jansen et al., 2022) and a Spearman correlation of 96% between SimilarWeb’s and NetMonitor’s ranks (Prantl and Prantl, 2018).
@@ 3.1. Month-over-month variation
    there is a seasonal pattern (similar to the one in (Liu et al., 2018)) in which traffic typically increases from September to October.
@@ 3.4. Manual inspection of top domains
We present our annotations in Tables 2 and 3 (see Appendix A). While the analysis of this subset of domains does not guarantee a comprehensive representation of total Web traffic, the trends observed within it are likely to extend to the following several thousand domains, thus reflecting the characteristics of a significant portion of Web usage.
@@ 4. Summary and conclusions
Additionally, we observe that a staggering 80% of the traffic to the top domains flows to companies in the United States. This substantial concentration, which could potentially characterize the Web as predominantly an American enterprise, has been previously highlighted in terms of companies’ market value (Evans and Gawer, 2016), but not in terms of traffic distribution. As discussed in Section 3.4, the accumulation of traffic on specific companies necessitates a socio-economic and political explanation. Hermes et al. took a step in this direction by interviewing European experts and top managers to glean insights into the reasons behind the dominance of American platforms in terms of market value (Hermes et al., 2020a). Their findings suggest that factors such as a results-oriented mindset, willingness to take risks, a sizable domestic market, state investment, early-mover advantage, the establishment of technology hubs (such as Silicon Valley), close collaborations with universities, and access to venture capital collectively contribute to American dominance in the web landscape.

### v2

@@ 1. Introduction
The importance of the Web, its lack of explicit management, and its vast potential have raised questions on what are its realized, actual characteristics in terms of main uses, applications, and content types, how visits are distributed among websites, who are the content publishers, how is the content typically produced, and much more. Answering these questions can provide valuable insights for developing public policies related to the Web, such as regulating new activities, relationships, and business models to promote positive outcomes and reduce negative impacts on society.
In the past, the Web has been thoroughly analyzed in terms of its content and graph properties (e.g., the number of webpages covering a given topic (Chakrabarti et al., 2002; Huizingh, 2000) and the number of hyperlinks connecting them (Adamic and Huberman, 2000b; Huberman, 2001)). However, previous accounts about the characteristics of the Web have primarily been based on metrics other than usage (e.g., link structure, number of websites, or companies’ market value) or were built on anecdotal accounts and non-quantitative data. Some exceptions are: (Adamic and Huberman, 2000a; Webster and Lin, 2002) that showed for a restricted subset of users and websites that visitors are distributed among websites in a power-law-like fashion; and (Agarwal and Sastry, 2022), that used Alexa’s ranking list to select the 100 most visited domains, Google trends to estimate their popularity and Fortiguard’s classification to verify the popularity of different topics, among other things.
@@
Our complete analysis can be accessed at http://github.com/cewebbr/web-unpacked.
@@ 2. Data Set
SimilarWeb, a company that combines all three methods, reportedly tracks a representative panel of users to ensure that all domains visited a minimum number of times are monitored (SimilarWeb, 2024b). Other companies that provide similar data are Semrush and Ahrefs.
@@
The quality of SimilarWeb data has been evaluated in previous academic works and is considered good enough for relative ballpark measurements between domains. Jansen et al. used Google Analytics for 86 websites as a truth table and identified a systematic bias of 20% in SimilarWeb’s monthly visits estimates (Jansen et al., 2022). Prantl compared monthly visits from SimilarWeb and NetMonitor [^1] https://www.spir.cz/projekty/netmonitor/ for 485 Czech websites and found an average absolute difference of about 42% between these two sets of measurements (Prantl and Prantl, 2018). Enterprise reports have indicated similar scenarios (Hardwick, 2018; Diachuk et al., 2021). However, given that our data on monthly visits cover over seven orders of magnitude (from 5 thousand to 86 billion), the reported deviations do not significantly affect the overall distribution of visits among websites. The data still provides precise ranks, with a Pearson correlation of 95% between monthly visit estimates from SimilarWeb and Google Analytics (Jansen et al., 2022) and a Spearman correlation of 96% between SimilarWeb’s and NetMonitor’s ranks (Prantl and Prantl, 2018). Also, the present work relies only on relative measurements, so any systematic biases are irrelevant.
It is important to note that web traffic exhibits annual patterns that could impact the representativeness of our three-month data sample (Liu et al., 2017, 2018). By analyzing five years of Google Trends data for the most visited domain names across various industries, we found that Western e-commerce sites like amazon.com and walmart.com see increased searches before Christmas, while searches for weather-related websites like weather.com and accuweather.com rise during school vacations. Assuming these annual patterns in Google Trends reflect similar patterns in actual domain visits, we estimate that our data may deviate from its annual average by up to 24%, as seen in the case of espn.com. However, major domains like youtube.com, instagram.com, and xvideos.com show only minor variations (i.e., less than 5%). These estimates suggest that, while seasonality is a factor, it is unlikely to affect our findings significantly.
@@ 3.1. Month-over-month variation
    there is a seasonal pattern (similar to the ones in (Liu et al., 2017, 2018)) in which traffic typically increases from September to October.
@@ 3.4. Manual inspection of top domains
We present our annotations in Tables 2 and 3 (see Appendix A), and in our repository. [^2] https://github.com/cewebbr/web-unpacked/blob/main/data/cleaned/domains-annotated_v03.csv While the analysis of this subset of domains does not guarantee a comprehensive representation of total Web traffic, the trends observed within it are likely to extend to the following several thousand domains, thus reflecting the characteristics of a significant portion of Web usage.
@@ 4. Summary and conclusions
Additionally, we observe that a staggering 80% of the traffic to the top domains flows to companies in the United States. This substantial concentration, which could potentially characterize the Web as predominantly an American enterprise, has been previously highlighted in terms of companies’ market value (Evans and Gawer, 2016), but not in terms of traffic distribution. As discussed in Section 3.4, the accumulation of traffic on specific countries necessitates a socio-economic and political explanation. Hermes et al. took a step in this direction by interviewing European experts and top managers to glean insights into the reasons behind the dominance of American platforms in terms of market value (Hermes et al., 2020a). Their findings suggest that factors such as a results-oriented mindset, willingness to take risks, a sizable domestic market, state investment, early-mover advantage, the establishment of technology hubs (such as Silicon Valley), close collaborations with universities, and access to venture capital collectively contribute to American dominance in the web landscape.